A disk of radius 3m contains a total charge 480 n


1
Expert's answer
2021-05-18T17:29:57-0400
"\rho=\frac{q}{\pi r^2}\\\rho=\frac{450\cdot10^{-9}}{\pi (3)^2}=1.6\cdot10^{-8}\frac{C}{m^2}"